OpenStudy (anonymous):

I was helping one member today and I got the impression that he/she did not know the method of writing a repeated decimal in terms of the sum of a geometric series. Here is an explanation of that topics.

OpenStudy (anonymous):

I will denote a repeated fraction as .baaaaaaaa ... .. For example if b=345 and a =6758 then the repeated fraction will be .3456758675867586758.... Let n=the length of b, in the example above n=3. Let m= the length of a, in the example above m=4. Let us try to write .baaaaaaa... as a sum in terms of a geometirc series. \[ .baaaaaaaa =\frac b {10^n}+\frac a {10^{m+n}}+\frac a {10^{m^2+n}}+ \frac a {10^{m^4+n}}+ \cdots =\\ \frac b {10^n}+\frac a {10^{m+n}}\left(1+\frac 1 {10^{m}}+ \frac 1 {10^{m^2}}+ \cdots \right)=\\ \frac b {10^n}+\frac a {10^{m+n}}\left( \frac {1 } {1- \frac 1 {10^{m}}} \right)\\ \]

OpenStudy (anonymous):

See also http://en.wikipedia.org/wiki/Geometric_series#Repeating_decimals

OpenStudy (experimentx):

Oo ..  this is quite unintuitive